OpenStudy (anonymous):

What is the value of work done on an object when a 50–newton force moves it 15 meters in the same direction as the force? 0 joules 3.3 joules 7.5 × 102 joules –7.5 × 102 joules

OpenStudy (anonymous):

I will help u

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

the Answer IS c

OpenStudy (shamim):

Work done w=F*s=50*15joule=?